OpenStudy (anonymous):

Which expression is the exponential form of square root of 30? 30^1 30^1/2 1/square root of 30 30^2

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):



OpenStudy (jdoe0001):

\(\huge \bf a^{\frac{n}{m}} = \sqrt[m]{a^n}\)

OpenStudy (skullpatrol):

That has no exponent in it?

OpenStudy (anonymous):

30^1/2

OpenStudy (skullpatrol):

Correct

OpenStudy (anonymous):

thanks

OpenStudy (jdoe0001):

yw

OpenStudy (skullpatrol):

Thanks for trying :)

OpenStudy (skullpatrol):

$$30^\frac{1}{2}$$